OpenStudy (anonymous):

Suppose it is 40 degrees inside your refrigerator and 360 degrees inside your oven when the power goes out. The refrigerator temperature starts rising at the rate of 5 degrees per hour. the oven temperature starts dropping at the rate of 45 degrees per hour. How long will it take for the temperatures to be the same? What will the temperature be then?

OpenStudy (anonymous):

replace A by wow i guess who ever wrote this problem was ignorant of newton's law of cooling, because this is not how it works at all no matter we can solve what is written lets call the time \(t\) in hours, then the temperature of the fridge is \[40+5t\] and the temp of the stove is \[360-45t\] set them equal get \[40+5t=360-45t\] and solve for \(t\)

OpenStudy (anonymous):

thank you so much!!!!!!!!!

OpenStudy (anonymous):

yw, hope it is clear how to solve it

OpenStudy (anonymous):

how would you find what the temperature would be now ?

OpenStudy (anonymous):

\[40+5t=360-45t\] \[50t=320\] \[t=320\div 50=6.4\]

OpenStudy (anonymous):

that doesnt make sense to me because the answer must be 70 degrees and i have to set up how you would figure it out. i dont understand how you would set it up.